the perimeter of a rectangle is 22 inches. The difference between the length and width is 3 inches. find the width
Answer by friesr(113)

Lets call the length x, then the width will be x-3. The perimeter of a rectangle is 2L + 2W
---
2x + 2(x-3) = 22
---
2x + 2x - 6 = 22
---
4x - 6 = 22
---
4x = 28
---
x = 28/4 or 7
The lenght = 7
Width = 7-3 or 4
---
Check
7 + 7 + 4 + 4 = 22